OpenStudy (anonymous):

find the equation of parabola with vertex (2,-3) directrix x=-2

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what's the distance from (2,-3) to the line x = -2

OpenStudy (anonymous):

k i have the answer can u check if it is correct or not

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

\[(y-3)^{2} =8(x-2)\]

OpenStudy (anonymous):

no it is +3

jimthompson5910 (jim_thompson5910):

that's not correct

jimthompson5910 (jim_thompson5910):

the directrix of (y+3)^2 =8(x-2) is x = 0, not x = -2

OpenStudy (anonymous):

but i got 8

OpenStudy (anonymous):

what is incorrect?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

it should be (y+3)^2 =16(x-2)

jimthompson5910 (jim_thompson5910):

the focal distance is p = 4 you would then multiply that by 4 to get 16

OpenStudy (anonymous):

opps i did 4*2 okay sorry

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

so the answer is (y+3)^2 = 16(x-2)

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

k i have 1 more

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

find the equation of hyperbola with vertices (2,-3) and (2,5) and foci (2,-4) and( 2,6)

jimthompson5910 (jim_thompson5910):

you got what

OpenStudy (anonymous):

\[\frac{ (x-2) ^{2}}{ 9 }- \frac{ (y-1)^{2} }{ 16 }\]

jimthompson5910 (jim_thompson5910):

that's not correct

OpenStudy (anonymous):

:(

jimthompson5910 (jim_thompson5910):

you're on the right track though

jimthompson5910 (jim_thompson5910):

you just have a few things mixed up

OpenStudy (anonymous):

like what?

jimthompson5910 (jim_thompson5910):

well the signs are mixed up for starters

jimthompson5910 (jim_thompson5910):

it should be in the form - ((x - h)^2)/(a^2) + ((y - k)^2)/(b^2) =1

OpenStudy (anonymous):

and what is the correct answer?

jimthompson5910 (jim_thompson5910):

you tell me, try it again and tell me what you get

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\frac{ (y-1) ^{2}}{ 16}-\frac{ (x-2) }{ 9 }=1\]

jimthompson5910 (jim_thompson5910):

much better

OpenStudy (anonymous):

so this is correct?

jimthompson5910 (jim_thompson5910):

oh you're missing a square over the (x-2), but yes it is

OpenStudy (anonymous):

oh sorry k thank you like i told u yeasterday day u are a great teacher

jimthompson5910 (jim_thompson5910):

thx and it's ok

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

yw